OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that contains the points (4, -7) and (0, 5).

OpenStudy (anonymous):

subtract y2 from y1 and divide the answer by x2-x1 [-3] y = -3x + c so far insert values of x and y in coordinate [0, 5] into equation to find c 5 =-3[0]+c 8 = 0 + c c= 8 hence answer is y = 8 -3x

OpenStudy (anonymous):

Error alert!!! [well it is late in France!] subtract y2 from y1 and divide the answer by x2-x1 [-3] y = -3x + c so far insert values of x and y in coordinate [0, 5] into equation to find c 5 =-3[0]+c 5 = 0 + c c= 5 hence answer is y = 5 -3x

OpenStudy (anonymous):

slope-intercept should look like: y=-3x+5 this is just a matter or presenting the answer. your solution is correct!